# Burst-tree structure and higher-order temporal correlations

### v1

@@ IV Results
To demonstrate the effects of kernels on higher-order temporal correlations, we visualize in Fig. 2 the entire event sequences of $n=300$ events generated using five kernels in Eqs. (1)–(5). For emphasizing the role of burst tree structure, we use exactly the same set of 299 interevent times (IETs) drawn from $P(\tau)$ in Eq. (6) with $\alpha=3.5$ and $\tau_{\rm c}=10^{7}$ for every event sequence. As expected, the event sequence using $K_{\rm const}$ shows no higher-order temporal correlations, compared to other event sequences showing such correlations. In the event sequence using $K_{\rm add}$, which is expected to lead to the heavy-tailed burst size distribution, one can observe bigger bursts than those in the case with $K_{\rm const}$, such as a burst in the interval of $270\lesssim t\lesssim 350$. The event sequence using $K_{\rm prod}$ obviously shows a very big burst corresponding to the gel formed in the coagulation process. Finally, two event sequences using kernels inspired by empirical results, namely, $K_{\rm emp1}$ and $K_{\rm emp2}$, show not only bigger bursts than those in the case with $K_{\rm const}$, but also the assortative mixing patterns of bursts which are not expected in the cases of constant, additive, and product kernels. The assortative mixing of bursts is evidenced, e.g., by bursts in the intervals of $10\lesssim t\lesssim 50$ and $80\lesssim t\lesssim 130$ for the case with $K_{\rm emp1}$, where relatively small bursts are followed by similarly small bursts.
@@
For more systematic comparison, we perform large-scale numerical simulations by generating $50$ different event sequences of $n=5\cdot 10^{5}$ events for each combination of five kernels in Eqs. (1)–(5) and the IET exponents $\alpha=1.8$, $2.5$, and $3.5$ of the IET distribution in Eq. (6). In all cases we set $\tau_{\rm c}=10^{7}$ in Eq. (6). Using the generated event sequences we measure the burst size distributions $Q_{s}(b)$ and $Q_{\Delta t}(b)$ for several values of $s$ and $\Delta t$, respectively, the memory coefficients for bursts $M_{\Delta t}$ for a wide range of $\Delta t$ in Eq. (8), and the autocorrelation function (ACF) $A(t_{\rm d})$ in Eq. (9). The results are summarized in Fig. 3.
Let us first look at the cases with constant, additive, and product kernels as shown in Fig. 3(a–l). The burst size distributions at several auxiliary time steps $s$, $Q_{s}(b)$, are exponential for the constant kernel as expected. For the additive and product kernels, $Q_{s}(b)$s at several values of $s$ show power-law scaling regimes with exponents $3/2$ and $5/2$, respectively. Burst size distributions $Q_{\Delta t}(b)$ obtained from the event sequences show essentially the same behaviors as $Q_{s}(b)$. We note that burst size distributions for the product kernel contain small peaks at the very large $b$ of the order of $n$, which are not shown in Fig. 3(i, j). These very large bursts correspond to the gels in the coagulation process.
@@
As mentioned, all of constant, additive, and product kernels may not lead to the assortative mixing of bursts, implying no correlations between consecutive burst sizes. Indeed, we observe $M_{\Delta t}\approx 0$ for these kernels. However, in the case of the constant kernel, we find $M_{\Delta t}\approx 0.2$ for small values of $\Delta t$ [Fig. 3(c)]. To explain this observation, we remind that for small $\Delta t$, there are a number of bursts of sizes $1$ or $2$. If these bursts were located in the sequence of burst sizes in a random manner, $M_{\Delta t}$ would be negligible. Instead, bursts of size $2$ seem to be located next to each other more often than expected, possibly due to the rule of randomly assigning left and right children nodes to the merged bursts. For example, if four bursts of sizes $1$, $1$, $2$, and $2$ are merged to two bursts of size $3$, which in turn are merged to a burst of size $6$, the sequence of burst sizes of $\{1,2,2,1\}$ may happen with a probability of $1/4$, eventually resulting in the positive $M_{\Delta t}$. This general argument however does not apply to the cases with additive and product kernels because the preferential mixing of bursts makes consecutive burst sizes differ from each other very quickly, leading to the negligible values of $M_{\Delta t}$. The negative values of $M_{\Delta t}$ for a very large $\Delta t$ are due to the finite-size effects as a few very big bursts and other relatively small bursts remain for such a large $\Delta t$. We finally remark that the IET exponent $\alpha$ affects only the timescales of the burst tree $\mathcal{T}$ not the ordinal burst tree $\mathcal{G}$; the smaller value of $\alpha$ implies a wider range of IETs and timescales. Therefore, the curve of $M_{\Delta t}$ for smaller $\alpha$ can be seen as a stretched version of that for larger $\alpha$.
@@
Table 1: Estimated values and fitting ranges of the decaying exponent $\gamma$ of ACFs presented in Fig. 3(d, h, l, p, t) for several values of the IET exponent $\alpha$ per kernel used. Values of the burst size exponent $\beta$ are exact for the constant, additive, and product kernels, while for the other kernels, values of $\beta$ are estimated from the numerical results of $Q_{\Delta t=64}(b)$ for $e^{2}\leq b\leq e^{7.5}$ in Fig. 3(n, r). Numbers in parentheses are errors.
Kernel $\beta$ $\alpha$ $\gamma$ Fitting range of $t_{\rm d}$
$K_{\rm const}(b,b^{\prime})$infty 1.8 0.28(1) [$e^{2}$, $e^{8}$]
@@
$K_{\rm emp1}(b,b^{\prime})$ 1.02(2) 1.8 0.18(1) [$e^{3}$, $e^{8}$]
2.5 0.22(1) [$e^{3}$, $e^{8}$]
3.5 0.39(2) [$e^{3}$, $e^{7}$]
$K_{\rm emp2}(b,b^{\prime})$ 1.28(2) 1.8 0.14(1) [$e^{2}$, $e^{10}$]
@@
In Fig. 3(d, h, l), we find that the ACFs decay algebraically with the time lag $t_{\rm d}$. The estimated values of the decay exponent $\gamma$ are summarized in Table 1. The event sequences generated using the constant kernel can be considered as results of the renewal process, in which there is no correlations between IETs. In that case, $\gamma$ has been known to be a function of $\alpha$ only; precisely, $\gamma=2-\alpha$ for $21<\alpha<$ 2 and $\gamma=\alpha-2$ for $\alpha>2$ [41]. However, numerical studies have shown strong finite-size effects on the mentioned scaling relation [40]. Considering such finite-size effects, the estimated values of $\gamma$ in Table 1 are consistent with the analytical results of $\gamma$. In the case of the additive kernel, $\beta=3/2$ implies strong correlations between consecutive IETs, obviously leading to very different behaviors than the results of the constant kernel. Note that although the analytical relation between $\alpha$, $\beta$, and $\gamma$ has recently been derived in Ref. [30], the derivation is based on the model considering only one value of $\Delta t$, hence we do not expect the analytical results in Ref. [30] account for our numerical observations considering the entire range of $\Delta t$. This strongly calls for a more comprehensive theoretical approach.
@@
The ACFs for the product kernel are found to be constant for a wide range of $t_{\rm d}$ as shown in Fig. 3(l), indicating $\gamma\approx 0$. One can get a hint from the visualized event sequence in Fig. 2, in which a very big burst containing almost all events is formed at a small value of $\Delta t$. The similar flat behaviors of the ACF have been reported in a model study based on the self-exciting point process [42], where the events form very big bursts that are separated by very large IETs (see Fig. 7 in Ref. [42]).
Simulation results using $K_{\rm emp1}$ and $K_{\rm emp2}$ are presented in Fig. 3(m–t). As expected, these kernels implement both preferential and assortative mixing patterns of bursts, leading to heavy-tailed burst size distributions and positive correlations between consecutive burst sizes at the same time. The estimated value of the burst size exponent $\beta$ for the $K_{\rm emp1}$ is $\approx 1$, indicating a heavier tail than that with $\beta=3/2$ for the additive kernel. Although the denominator in Eq. (4) was introduced to implement the assortative mixing, it also seems to enhance the preferential mixing by effectively reducing the possibility of small bursts being chosen. On the other hand, the estimated value of $\beta\approx 1.28(2)$ for the $K_{\rm emp2}$ is more similar to $\beta=3/2$ for the additive kernel than to $5/2$ for the product kernel. That is, the term of $\ln b+\ln b^{\prime}$ in the first parenthesis in Eq. (5) can be seen being more similar to $b+b^{\prime}$ than $bb^{\prime}$. Interestingly, the estimated values of $\gamma$ for $\alpha=1.8$ and $2.5$ are also similar to those for the additive kernel, which is not the case with $\alpha=3.5$. Finally, as shown in Fig. 3(o, s), the memory coefficients for bursts are positive for a wide range of $\Delta t$ in both cases of $K_{\rm emp1}$ and $K_{\rm emp2}$. Thus these kernels generate most realistic event sequences as expected.
@@ V Conclusion
We have studied how different burst tree structures generated using several burst-merging kernels shape the higher-order temporal correlations in the time series. For this, we adopt constant, additive, and product kernels as well as two kernels inspired by the previous empirical results [20]. To characterize the higher-order temporal correlations, we measure the burst size distributions, memory coefficients for bursts, and the autocorrelation function. We numerically find that kernels with preferential mixing lead to the heavy-tailed burst size distributions, while kernels with assortative mixing lead to positive correlations between burst sizes. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the interevent time distribution. Our results in this work would shed lights on understanding various empirical results reported in Ref. [20].
As mentioned, thanks to the analogy to the coagulation process, analytical solutions of burst size distributions for the constant, additive, and product kernels could be obtained, while those for more complicated kernels are not straightforward to obtain. Analytical approaches to the analytical solutions for such more complicated kernels as well as to the relation between $\alpha$, $\beta$, and $\gamma$ in more general cases are left for future works.
@@ Acknowledgements.
T.B. and H.-H.J. acknowledge financial support by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A2C1007358). This work was supported by the Research Fund of The Catholic University of Korea in 2024.

### v2

@@ IV Numerical results
To demonstrate the effects of kernels on higher-order temporal correlations, we visualize in Fig. 2 the entire event sequences of $n=256$ events generated using seven kernels in Eqs. (1)–(7). For emphasizing the role of burst tree structure, we use exactly the same set of 255 interevent times (IETs) drawn from $P(\tau)$ in Eq. (8) with $\alpha=3.5$ and $\tau_{\rm c}=10^{7}$ for all cases. As expected, the event sequence using $K_{\rm const}$ shows no higher-order temporal correlations, compared to other event sequences showing such correlations. In the event sequence using $K_{\rm sum}$, which is expected to lead to the heavy-tailed burst size distribution, one can observe bigger bursts than those in the case with $K_{\rm const}$, such as bursts in the interval of $190\lesssim t\lesssim 260$. The event sequence using $K_{\rm prod}$ obviously shows a very big burst corresponding to the gel formed in the coagulation process. The other four event sequences are generated kernels having the feature of assortative merging; all of them show not only bigger bursts than those in the case with $K_{\rm const}$, but also bursts of similar sizes following each other, which are not expected in the cases of constant, sum, and product kernels. Such an assortative pattern is evidenced most clearly in the case with $K_{\rm diag}$. It is also seen for the case with $K_{{\rm sum\text{-}diag}}$, e.g., in the intervals of $10\lesssim t\lesssim 50$ and $80\lesssim t\lesssim 130$, where relatively small bursts are followed by similarly small bursts. Finally, we note that although $K_{\rm diag}$ does not implement the preferential merging by definition, the generated event sequence shows some bigger bursts than those by $K_{\rm const}$, possibly implying that the heavy-tailed burst size distributions may be observed due to the assortative merging only. This issue will be discussed later.
@@
For the systematic comparison, we perform large-scale numerical simulations by generating $50$ different event sequences for each combination of seven kernels in Eqs. (1)–(7) and the IET exponents $\alpha=1.8$, $2.5$, and $3.5$ of the IET distribution with $\tau_{\rm c}=10^{7}$ in Eq. (8). The number of events in each event sequence is $n=5\cdot 10^{5}$ for all kernels except for $K_{\rm diag}$ and $K_{{\rm diag\text{-}pre}}$, for which $n=2^{19}=524,288$ has been used. Using the generated event sequences we measure burst size distributions $Q_{\Delta t}(b)$ for several values of $\Delta t$, memory coefficients for bursts $M_{\Delta t}$ for a wide range of $\Delta t$ in Eq. (10), and the autocorrelation function (ACF) $A(t_{\rm d})$ in Eq. (11).
@@
Let us first look at the cases with constant, sum, and product kernels as shown in Fig. 3. The burst size distributions at several timescales, $Q_{\Delta t}(b)$, are exponential for $K_{\rm const}$ as expected. For $K_{\rm sum}$ and $K_{\rm prod}$, $Q_{\Delta t}(b)$s at several values of ${\Delta t}$ show power-law scaling regimes with exponents $3/2$ and $5/2$, respectively. Note that burst size distributions for $K_{\rm prod}$ contain small peaks at the very large $b$ of the order of $n$, which are not shown in Fig. 3(g). These very large bursts correspond to the gels in the coagulation process. As mentioned, all of constant, sum, and product kernels may not lead to positive correlations between consecutive burst sizes. Indeed, we observe $M_{\Delta t}\approx 0$ for these kernels. However, in the case of $K_{\rm const}$, we find $M_{\Delta t}\approx 0.2$ for small values of $\Delta t$ [Fig. 3(b)]. To explain this observation, we remind that for small $\Delta t$, there are a number of bursts of sizes $1$ or $2$. If these bursts were located in the sequence of burst sizes in a random manner, $M_{\Delta t}$ would be negligible. Instead, bursts of size $2$ seem to be located next to each other more often than expected, possibly due to the rule of randomly assigning left and right children nodes to the merged bursts. For example, if four bursts of sizes $1$, $1$, $2$, and $2$ are merged to two bursts of size $3$, which in turn are merged to a burst of size $6$, the sequence of burst sizes of $\{1,2,2,1\}$ may happen with a probability of $1/4$, eventually resulting in the positive $M_{\Delta t}$. This general argument however does not apply to the cases with $K_{\rm sum}$ and $K_{\rm prod}$ because the preferential merging forces consecutive burst sizes to differ from each other very quickly, leading to the negligible values of $M_{\Delta t}$. The negative values of $M_{\Delta t}$ for a very large $\Delta t$ are due to the finite-size effects as a few very big bursts and other relatively small bursts remain for such a large $\Delta t$. We finally remark that the IET exponent $\alpha$ affects only the timescales of the burst tree $\mathcal{T}$ not the ordinal burst tree $\mathcal{G}$; the smaller value of $\alpha$ implies a wider range of IETs and timescales. Therefore, the curve of $M_{\Delta t}$ for smaller $\alpha$ can be seen as a stretched version of that for larger $\alpha$.
@@
Table 1: Estimated values and fitting ranges of the decaying exponent $\gamma$ of ACFs presented in Figs. 3 and 4 for several values of the IET exponent $\alpha$ per kernel used. “P” and “A” mean the preferential and assortative merging rules, respectively. Values of the burst size exponent $\beta$ are exact for the constant, sum, and product kernels [41], while for the other kernels, values of $\beta$ are estimated from the numerical results of $Q_{\Delta t=64}(b)$. Numbers in parentheses are errors.
Kernel P A $\beta$ $\alpha$ $\gamma$ Fitting range of $t_{\rm d}$
$K_{\rm const}$$\times$ $\times$ $\infty$ 1.8 0.28(1) [$e^{2}$, $e^{8}$]
@@
$K_{\rm diag}$$\times$ 1.00(1) [^1] Fitted for $b\in[e^{2.5},e^{8.5}]$. 1.8 0.23(1) [$e^{3}$, $e^{12}$]
2.5 0.36(1) [$e^{2}$, $e^{9}$]
3.5 0.77(7) [$e^{2}$, $e^{8}$]
$K_{{\rm diag\text{-}pre}}$ 0.85(2) [^2] Fitted for $b\in[e^{2.5},e^{7}]$. 1.8 0.089(1) [$e^{3}$, $e^{9}$]
2.5 0.14(1) [$e^{3}$, $e^{8}$]
3.5 0.33(1) [$e^{3}$, $e^{6}$]
$K_{{\rm sum\text{-}diag}}$ 1.02(2) [^3] Fitted for $b\in[e^{2},e^{7.5}]$. 1.8 0.19(1) [$e^{2}$, $e^{7}$]
2.5 0.20(1) [$e^{3}$, $e^{8}$]
3.5 0.38(1) [$e^{2}$, $e^{7}$]
$K_{\rm emp}$ 1.28(2) [^4] Fitted for $b\in[e^{2},e^{7.5}]$. 1.8 0.14(1) [$e^{2}$, $e^{10}$]
@@
Figure 4: The same as Fig. 3 but using $K_{\rm diag}$, $K_{{\rm diag\text{-}pre}}$, $K_{{\rm sum\text{-}diag}}$ with $c_{1}=0.01$, and $K_{\rm emp}$ with $c_{2}=3$, $c_{3}=100$, and $c_{4}=0.25$ [Eqs. (4)–(7)] (from top to bottom). For $K_{\rm diag}$ and $K_{{\rm diag\text{-}pre}}$, we use $n=2^{19}=524,288$. Black lines in panels (a, d, g, j) show fitting ranges and results of the burst-size exponent $\beta$ in Eq. (9), also summarized in Table 1.
@@
In Fig. 3(c, f, i), we find that the ACFs decay algebraically with the time lag $t_{\rm d}$. The estimated values of the decay exponent $\gamma$ are summarized in Table 1. The event sequences generated using $K_{\rm const}$ can be considered as results of the renewal process, in which there is no correlations between IETs. In that case, $\gamma$ has been known to be a function of $\alpha$ only; precisely, $\gamma=2-\alpha$ for $21<\alpha<$ 2 and $\gamma=\alpha-2$ for $\alpha>2$ [43]. However, numerical studies have shown strong finite-size effects on the mentioned scaling relation [42]. Considering such finite-size effects, the estimated values of $\gamma$ in Table 1 seem to be overall consistent with the analytical results of $\gamma$. In the case of $K_{\rm sum}$, $\beta=3/2$ implies long-range temporal correlations between consecutive IETs, leading to different values of $\gamma$ than those for $K_{\rm const}$. Although there are no analytical solutions of the ACF for $K_{\rm sum}$, one can refer to the recent analytical result of $\gamma$ as a function of $\alpha$ and $\beta$ in Ref. [30]; for the range of $21<\beta<$ 2, $\gamma$ is an increasing function of $\alpha$. Thus, it is qualitatively consistent with the numerical values of $\gamma$ for $K_{\rm sum}$ in Table 1. We remark that since the model studied in Ref. [30] considers only one timescale $\Delta t$, its analytical results might systematically deviate from numerical results of our model considering the entire range of $\Delta t$.
@@
The ACFs for $K_{\rm prod}$ are found to be constant for a wide range of $t_{\rm d}$ as shown in Fig. 3(i), indicating $\gamma\approx 0$. One can get a hint from the visualized event sequence in Fig. 2, in which a very big burst containing almost all events is formed at a small value of $\Delta t$. The similar flat behaviors of the ACF have been reported in a model study based on the self-exciting point process [44], where the events form very big bursts that are separated by very large IETs (see Fig. 7 in Ref. [44]).
Simulation results using the other four kernels are presented in Fig. 4. We first discuss the case with the diagonal kernel $K_{\rm diag}$ in Eq. (4); although it does not have a feature for the preferential merging, the power-law regime with $\beta\approx 1$ is clearly observed in the burst size distributions. This indicates that the heavy-tailed burst size distributions might be not only due to the preferential merging but also due to the assortative merging. To understand this observation, one can map the merging process to a system of particles on one-dimensional lattice, where each site $i$ corresponds to the burst size $b=2^{i}$ ($i=0,1,\ldots$), all particles are initially at the site 00, and two particles at the same site $i$ hop to the next site $i+1$ after merging. Since the probability of choosing two particles is constant irrespective of their position, the number of particles at the site $i$ can be written as $n_{i}\propto c^{i}$ with a constant $c\in(0,1)$, which is combined with $b=2^{i}$ to give $Q(b)\propto b^{-\beta}$ with $\beta=-\log_{2}c>0$. We also find that $M_{\Delta t}>0$ for the entire range of $\Delta t$, and that $\gamma$ increases with $\alpha$, again qualitatively consistent with the analytical result in Ref. [30].
Finally, we find that the other kernels, i.e., $K_{{\rm diag\text{-}pre}}$, $K_{{\rm sum\text{-}diag}}$, and $K_{\rm emp}$, which implement both preferential and assortative merging, lead to heavy-tailed burst size distributions and positive correlations between consecutive burst sizes at the same time, as expected. From now on we focus on the scaling behaviors of $Q_{\Delta t}(b)\propto b^{-\beta}$ and $A(t_{\rm d})\propto t_{\rm d}^{-\gamma}$ because $M_{\Delta t}>0$ is naturally expected in all cases.
The estimated value of $\beta$ for $K_{{\rm diag\text{-}pre}}$ is $\approx 0.85$, indicating a heavier tail than that for $K_{\rm diag}$. This can be understood by considering the fact that $K_{{\rm diag\text{-}pre}}$ has a preferential factor $b$ multiplied to $K_{\rm diag}$. However, such a numerical result is not consistent with the analytical solution in Eq. (16), requiring further investigation in the future. Next, by comparing $\beta\approx 1$ for $K_{{\rm sum\text{-}diag}}$ to $\beta=3/2$ for $K_{\rm sum}$, we observe that the tails of burst size distributions get heavier due to the assortative merging with $c_{1}>0$ in Eq. (6). We also find $\beta\approx 1.28$ for $K_{\rm emp}$. It turns out that estimated values of $\beta$ for these three kernels are smaller than $2$, allowing us to expect that $\gamma$ increases with $\alpha$ [30], which is indeed the case as shown in Table 1. The scaling behaviors for the case with $K_{\rm emp}$ appear most similar to those with $K_{\rm sum}$, except for the value of $\gamma$ when $\alpha=3.5$ is used.
@@ V Conclusion
We have studied how different burst-merging kernels shape the higher-order temporal correlations in the burst tree structure as well as the time series. For this, we adopt constant, sum, product, and diagonal kernels as well as those inspired by the previous empirical results [20]. To characterize the higher-order temporal correlations, we measure the burst size distributions, memory coefficients for bursts, and the autocorrelation function. By large-scale numerical simulations, we generically find that kernels with preferential merging lead to the heavy-tailed burst size distributions, while kernels with assortative merging lead to positive correlations between burst sizes. We also find that in some kernels the assortative merging enhances the heavy tail in burst size distributions. The decaying exponent of the autocorrelation function depends not only on the kernel but also on the power-law exponent of the interevent time distribution. Therefore, our findings would shed lights on understanding various empirical results reported in Ref. [20].
Despite the usefulness of our framework using burst-merging kernels in generating or analyzing higher-order temporal correlations, underlying mechanisms behind such kernels are not yet fully understood. The kernels are used to determine which bursts are merged to become bigger bursts, rather than explain the reasons for such a determination at the fundamental level. Thus, it would be interesting if one could devise more mechanistic processes, e.g., based on queuing models [8, 45, 46] or self-exciting point processes [47, 48, 49, 44], and translate them into the burst-merging kernels. This topic is left for future works.
@@ Acknowledgements.
H.-H.J. thanks Mikko Kivelä and Takayuki Hiraoka for fruitful discussions. T.B. and H.-H.J. acknowledge financial support by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A2C1007358). This work was supported by the Research Fund of The Catholic University of Korea in 2024.